A tanker is travelling with a velocity of v = [30, 15] and a forklift on board is traveling relative to the tanker with a velocityof v =–9ir +7rj.A spider is crawling on the forklift with velocity v = – i + 5 k relative to the forklift. Calculate the velocity and speed of the spider relative to the ocean.
Answer #1

The velocity of the tanker relative to ocean = 30 i + 15 j

The velocity of forklift with respect to the tanker = -9 i + 7 j

The velocity of spider with respect to forklift = -i + 5k

So,velocity of spider with respect to tanker = (-i + 5k) + (-9 i + 7 j) = -10 i + 7 j + 5 k

The velocity of the spider with respect to the ocean = (-10 i + 7 j + 5 k) + (30 i + 15 j) = 20 i +22 j + 5 k

The speed is given by

speed = \sqrt{20^2+22^2+5^2} = 30.149 m/s